OpenStudy (anonymous):

a painter leans a 20 foot ladder

geerky42 (geerky42):

Cool!

OpenStudy (anonymous):

lmao @geerky42 @samanthasilva what is the rest of the question.?

OpenStudy (anonymous):

From the problem description, when the painter releases the ladder, it will fall to the ground.

OpenStudy (anonymous):

a^2+b^2=c^2 a=16, c=20 256+b^2=400 b^2=144 b=12' dist of ladder from wall.

OpenStudy (anonymous):

base=12ft, height=16ft ladder=20ft b^2=c^2-a^2 b^2=20^2-12^2 b^2=400-144 b^2=256 b=sq rt of 256 =16ft (height)